OpenStudy (anonymous):

Solve the following system by graphing. x - y = 4 x + y = 2 What is the solution of the system? (3, -1) (3, 1) (-1, 3)

OpenStudy (lgbasallote):

have you tried graphing?

OpenStudy (anonymous):

GRAPH!!

OpenStudy (anonymous):

Add the two equations and you get x - y = 4 x + y = 2 2 x = 6 x =3

OpenStudy (anonymous):

Solve the following system by graphing.

OpenStudy (anonymous):

Find y by replacing in one of the equations x - y = 4 3 -y =4 y=-1